# Development of Neural Network-Based Optimal Control Pulse Generator for Quantum Logic Gates Using the GRAPE Algorithm in NMR Quantum Computer

### v1

@@ Abstract
In this paper, we introduce a neural network to generate the optimal control pulses for general single-qubit quantum logic gates, within a Nuclear Magnetic Resonance (NMR) quantum computer. By utilizing a neural network, we can efficiently implement any single-qubit quantum logic gates within a reasonable time scale. The network is trained by control pulses generated by the GRAPE algorithm, all starting from the same initial point. After implementing the network, we tested it using numerical simulations. Also, we present the results of applying Neural Network-generated pulses to a three-qubit benchtop NMR system and compare them with simulation outcomes. These numerical and experimental results showcase the precision of the Neural Network-generated pulses in executing the desired dynamics. Ultimately, by developing the neural network using the GRAPE algorithm, we discover the function that maps any single-qubit gate to its corresponding pulse shape. This model enables the real-time generation of arbitrary single-qubit pulses. When combined with the GRAPE-generated pulse for the CNOT gate, it creates a comprehensive and effective set of universal gates. This set can efficiently implement any algorithm in noisy intermediate-scale quantum computers (NISQ era), thereby enhancing the capabilities of quantum optimal control in this domain. Additionally, this approach can be extended to other quantum computer platforms with similar Hamiltonians.
@@ I Introduction
There are several factors that limit the fidelity of quantum logic gates, including the imprecision of classical control instruments, interactions of qubits with the environment, and material imperfections and manufacturing defects in the quantum hardware layer. Most of these factors are related to the platform used to set up the quantum processor. For instance, cross-talk between qubits and the inevitable evolution of qubits due to qubit-mediated interactions during single-qubit gates are significant issues. These challenges arise even in closed quantum systems and are prevalent across various platforms, including NMR and superconducting systems. This is where optimal quantum control methods can play a crucial role in enhancing gate fidelities.
@@
The NMR platform is an excellent choice for testing quantum control ideas. Firstly, bench-top NMR spectrometers are now accessible in many university laboratories, making them convenient for research. Additionally, due to their weak magnetic field, advanced pulse shaping is essential for optimal control in bench-top NMR quantum computers. Secondly, the results obtained can be readily applied and extended to other platforms with similar Hamiltonians, such as controlling superconducting qubits in a quantum bus resonator.
@@
Quantum logic gates are implemented on the NMR platform by applying RF magnetic field pulses, where, the central frequencies of these pulses close to the Larmor frequencies of different nuclei [1, 2, 3]. Extensive research has been conducted to determine the various pulse shapes required to implement desired operations with high fidelity [4, 5, 6, 7, 8]. One method for finding the optimal pulse shapes for specific logic gates involves using optimal control theory, which offers several approaches [9, 10, 11, 12, 13]. A notable numerical approach within this theory is the GRAPE algorithm [14]. This method focuses on finding the optimal pulse shapes for specific quantum logic gates (e.g., S, T, H, X, Y, and Z) [15]. Once the pulse shapes for these basic gates are generated, which can take a relatively long time, they are registered in the compiler. These registered pulses are then used to implement any other arbitrary gates in the algorithm.
@@
A more sophisticated goal, particularly crucial in the NISQ era, is to extend this approach to design pulse shapes for directly implementing arbitrary gates without decomposing them into basic gates (i.e., implementing any arbitrary single-qubit gate with a single pulse) [16]. This is where neural networks come into play. Now, the network itself can be considered part of the compiler, allowing for the generation of pulses for any arbitrary gates in a reasonable time, i.e., compile time.
In recent years, there has been a surge in research exploring the application of machine learning algorithms within the realm of quantum optimal control theory [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Also, there has been significant research in developing NMR or MRI RF control pulse designs using neural networks by utilizing the excitation profile as a training set [30, 31, 32, 33, 34, 35]. Furthermore, the usage of neural networks in the shimming of NMR spectrometer is discussed [36, 37, 38]. In this work, we have developed the neural network, by applying the GRAPE algorithm for sampling to train networks for implementing the logical gates in NMR systems. The concept of developing control pulses can be expanded through the use of neural networks, as explained here [16, 39]. We used and extended the main ideas presented in [16, 39], by developing the method of sampling to train the neural network and testing it in an NMR quantum computing implementation. Utilizing a neural network, we can input the desired operation and, in return, receive the pulse shape as the output. This provides us with a versatile tool capable of generating pulse shapes for any arbitrary gates in a reasonable time, rather than being limited to a single specific gate, which is the limitation of the GRAPE algorithm [16, 39].
@@
To design such a network, a training set is required. For efficient training, we focus on generating phase-only control pulses [40, 41]. One straightforward strategy involves generating a series of random pulses and computing the resulting operations they enact on the system. Consequently, we obtain numerous ordered pairs: the first element is the operation, and the second is the corresponding pulse shape. With this training set, we can proceed to train a neural network. However, when we applied this approach to a three-qubit NMR system, the results were sub-optimal, and the error rates failed to decrease satisfactorily. We then considered utilizing the GRAPE algorithm to develop a training set for the neural network. This time, we selected a variety of single-qubit gates at random and employed the GRAPE algorithm to determine the optimal control pulses for their implementation. A critical aspect of using the GRAPE algorithm for developing these pulses was initiating the search from the same starting point to find the optimal results. After calculating the cosine similarity between the randomly generated pulses in the previous approach, we found that the similarity was not particularly significant. By beginning from a same point in the GRAPE algorithm to develop pulses for the desired gates, we effectively narrowed down a specific region in the space of control parameters. This means that within this region, the pulses are closely related to one another, Fig.1.
@@
As we will demonstrate, the cosine similarity between the resulting pulses has significantly improved. This approach helps us to focus on data around a specific point in the solution space. Intuitively, starting from a random point in each GRAPE iteration causes the algorithm to converge to different minima with the same fidelity value. In a problem with multiple, or even infinitely many, local minima, it is crucial to limit our analysis of the dynamic system to a specific local solution. Ignoring this necessity results in severely non-smooth data geometry, preventing a simple, feasible neural network from finding any patterns within the feature space.
@@
We now possess a neural network that, upon inputting your desired single-qubit gate, outputs the shape of the pulse. Another achievement of this neural network approach is its ability to implement complicated operations in a single pulse. This method is advantageous over executing gates sequentially, particularly in the NISQ era, where the fidelities of individual gates would compound, resulting in a reduced overall fidelity. By directly generating the operation, we can achieve it with significantly higher fidelity and fewer steps.
@@ II Hamiltonian of the system and form of the control pulses
For the n-qubit NMR system with homo-nuclear molecule, one can demonstrate that if we are in the liquid regime and at room temperature, and if we have only one RF channel, then by choosing a rotating frame which rotates with the RF frequency of the channel, we can derive the following terms for the total Hamiltonian [42, 2, 3, 4]:
@@
$H_{\text{ctrl}}=\omega_{1}(t)(\cos(\phi(t))\sum_{i=1}^{n}\sigma_{x}^{(i)}+\sin(\phi(t))\sum_{i=1}^{n}\sigma_{y}^{(i)}),$ (2)
$\gamma_{i}B_{1}(t)=\omega_{1}(t),\gamma_{i}B_{0}=\omega_{0,i},$ (3)
where $B_{0}$ is the static magnetic field, and $B_{1}(t)$ and $\phi(t)$ represent the amplitude and the phase of the applied RF field, respectively. $\gamma_{i}$ is the gyromagnetic ratio for each nucleus in the system. Because $B_{1}(t)$ is significantly smaller than $B_{0}$, and the system is homo-nuclear we used the approximation $\gamma_{i}B_{1}(t)=\omega_{1}(t)$ for any $i$. The $H_{\text{ctrl}}$ and the first term in $H_{\text{free}}$ are due to the Zeeman interaction and the second term in $H_{\text{free}}$ represents the J-coupling between different nuclei [3].
To implement the desired quantum logic gates, we must determine $B_{1}(t)$ and $\phi(t)$. As a common technique in the GRAPE algorithm, these pulses are approximated as constant within $N$ discrete time steps. Within each step, the amplitude and phase remain constant, allowing us to compute the system’s overall operation by multiplying the operations from each time step. Since the Hamiltonian is constant within each step, we can readily calculate the dynamics for that period. Consequently, the total operation resulting from the pulse sequence is obtained by the product of the dynamics across all steps [43]. In this algorithm, the derivative of fidelity with respect to control parameters is computed. By iteratively updating these parameters and moving in the direction of the gradient, we can converge on the optimal solution [15]. To simplify the learning process and reduce the number of control parameters, we used only phases to develop control pulses while keeping the amplitude constant. As we will see (and discussed in [40, 41]), this constraint does not significantly limit us in achieving acceptable fidelities, and we can reach high fidelities by only modulating the phase. To test the stability of this approach, after calculating the phase-modulated pulse with constant amplitude, we used it as the initial state for the GRAPE algorithm, by adding the amplitudes as an additional control parameters. The final result was not significantly different from the initial state, indicating that these pulses approximately satisfy the condition of zero gradient Euclidean norm with respect to phase elements, even after additional degrees of freedom related to amplitudes are added to the optimization problem.
@@ III Designing the neural network
As previously discussed, efficiently training the target neural network requires a set of samples with high cosine similarity. One straightforward strategy involves generating a series of random pulses and computing the resulting operations they enact on the system. After calculating the cosine similarity between the randomly generated pulses in this approach, we found that the similarity was not particularly significant.
We then considered utilizing the GRAPE algorithm to develop a training set for the neural network. This time, we selected a variety of single-qubit gates at random and used the GRAPE algorithm to determine the optimal control pulses for their implementation, each with the same starting point. Additionally, we examined $16$ samples generated using the GRAPE algorithm for a specific logic gate, each with a different starting point.
A comparison between Fig.2 and Fig.3 illustrates a noticeable increase in similarity among the samples when generated using the GRAPE algorithm with the same starting point. Therefore, a critical aspect of using the GRAPE algorithm for generating the training pulses is initiating the search from the same starting point. By beginning from the same point in the GRAPE algorithm to seek pulses for the desired gates, we effectively narrowed down a specific region in the space of control parameters. This means that within this region, the pulses are closely related to one another.
Finally, we selected the random unitary operations and subsequently used the GRAPE algorithm, initiating from the same starting point, to generate the control pulses necessary for their implementation. After creating 17,000 samples, we trained a feed-forward neural network [44] to establish the mapping function between the the overall operation applied to the system and control pulses. To implement the layers of the neural network, we utilized the architecture depicted in Fig. 5.
@@
Figure 2: Cosine similarity analysis of 16 samples generated using the GRAPE algorithm for a specific logic gate, each with a different starting point. Value of the similarity between the pulses x and y is shown on the x-y cells.
@@
Figure 3: Cosine similarity analysis of 17,000 samples generated using the GRAPE algorithm for the random logic gates, each with a same starting point. Value of the similarity between the pulses x and y is shown on the x-y cells.
Figure 4: The neural network identifies the operation on the input and predicts the pulse required to implement the desired gate, which is then transmitted to the output. a, b, c and d are complex numbers.
Figure 5: Layers of the neural network include: fc, which stands for fully connected network; ReLU and tanh, which are related to activation functions; and dropout, which is used to prevent the network from overfitting the data.

### v2

@@ Abstract
In this paper, we introduce a neural network to generate optimal control pulses for general single-qubit quantum logic gates, within a Nuclear Magnetic Resonance (NMR) quantum computer. By utilizing a neural network, we can efficiently implement any single-qubit quantum logic gates within a reasonable time scale. The network is trained by control pulses generated by the GRAPE algorithm, all starting from the same initial point. After implementing the network, we tested it using numerical simulations. Also, we present the results of applying Neural Network-generated pulses to a three-qubit benchtop NMR system and compare them with simulation outcomes. These numerical and experimental results showcase the precision of the Neural Network-generated pulses in executing the desired dynamics. Ultimately, by developing the neural network using the GRAPE algorithm, we discover the function that maps any single-qubit gate to its corresponding pulse shape. This model enables the real-time generation of arbitrary single-qubit pulses. When combined with the GRAPE-generated pulse for the CNOT gate, it creates a comprehensive and effective set of universal gates. This set can efficiently implement any algorithm in noisy intermediate-scale quantum computers (NISQ era), thereby enhancing the capabilities of quantum optimal control in this domain. Additionally, this approach can be extended to other quantum computer platforms with similar Hamiltonians.
@@ I Introduction
Several factors limit the fidelity of quantum logic gates, including the imprecision of classical control instruments, manufacturing defects in the quantum hardware layer, and qubit interactions with the environment, such as material imperfections. Additionally, some challenges arise from unwanted interactions within the quantum layer itself, even in closed quantum systems. For example, qubit cross-talk and the inevitable evolution of qubits due to qubit-mediated interactions during single-qubit operations present significant obstacles. These issues are prevalent across various platforms, including NMR and superconducting systems. This is where optimal quantum control methods can play a crucial role in enhancing gate fidelities.
In multi-qubit quantum computing systems, achieving high-fidelity gate operations poses a significant quantum control challenge. In our work, this involves employing pulse shaping techniques to precisely guide the target qubit through its intended evolution while ensuring that non-target qubits remain effectively undisturbed. However, inherent system couplings often induce unwanted interactions, leading to erroneous evolution of the target qubit and spurious dynamics in non-target qubits. To mitigate these effects, we engineer control pulses that direct the target qubit through its intended evolution while simultaneously steering non-target qubits along compensatory trajectories, such that their quantum states are restored by the conclusion of the target qubit’s operation. This strategy maintains the locality of the transformation, suppresses cross-talk, and enhances the overall fidelity of the control scheme by confining the unitary evolution to the intended subsystem.
@@
The NMR platform is an excellent choice for testing quantum control ideas. First, bench-top NMR spectrometers are widely accessible in university laboratories, making them convenient for research. Moreover, due to their weak magnetic field, advanced pulse shaping is crucial for achieving optimal control in bench-top NMR quantum computers. Additionally, the results obtained can be readily applied and extended to other platforms with similar Hamiltonians, such as the control of superconducting qubits in a quantum bus resonator.
@@
Quantum logic gates are implemented on the NMR platform by applying RF magnetic field pulses, with their central frequencies positioned close to the Larmor frequencies of different nuclei [1, 2, 3]. Extensive research has been conducted to determine the optimal pulse shapes required for high-fidelity quantum operations [4, 5, 6, 7, 8]. One approach to finding these pulse shapes for specific logic gates is optimal control theory, which offers several methodologies [9, 10, 11, 12, 13]. A particularly notable numerical technique within this framework is the GRAPE algorithm [14]. This method is designed to optimize pulse shapes for basic quantum logic gates, such as S (phase gate), T (${\pi}/{8}$ gate), H (Hadamard gate), and the Pauli gates X, Y, and Z [15]. Once the pulse shapes for these fundamental gates are generated—a process that is time-intensive—they are registered in the compiler. These registered pulses are then used to efficiently implement arbitrary gates within the algorithm.
@@
A more sophisticated goal, particularly crucial in the NISQ era, is to extend this approach to designing pulse shapes for directly implementing arbitrary gates without decomposing them into basic ones—essentially enabling the execution of any arbitrary single-qubit gate with a single pulse [16]. This is where neural networks come into play. By integrating the network into the compiler, pulse generation for arbitrary gates can be achieved within a reasonable time, i.e., during compile time.
In recent years, there has been a surge in research exploring the application of machine learning algorithms within the realm of quantum optimal control theory [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29]. Also, there has been significant research in developing NMR or MRI RF control pulse designs using neural networks by utilizing the excitation profile as a training set [30, 31, 32, 33, 34, 35]. Furthermore, neural networks have been explored for their role in the shimming of NMR spectrometers [36, 37, 38], a feedback control process that ensures a homogeneous magnetic field across the NMR sample via various magnet coils. This process allows identical nuclei to experience uniform Zeeman splitting and have identical Larmor frequencies, resulting in narrower spectral lines.
In this work, we develop a neural network by employing the GRAPE algorithm for sampling, enabling the training of networks to implement logical gates in NMR quantum computers. The concept of control pulse generation can be further expanded through neural networks, as discussed in [16, 39]. Building on the foundational ideas presented in [16, 39], we enhance the sampling method for training the neural network and test its effectiveness in an NMR quantum computing implementation. By leveraging a neural network, we can input the desired operation and obtain the corresponding pulse shape as output. This approach provides a versatile tool capable of generating pulse shapes for arbitrary gates in a reasonable time scale, overcoming the limitation of the GRAPE algorithm, which is constrained to optimizing pulse shapes for individual specific gates [16, 39].
To design such a network, a well-structured training set is required. For efficient training, we focus on generating phase-only control pulses [40, 41]. A straightforward strategy involves generating a series of random pulses and obtaining their resulting operations on the system. This process yields numerous ordered pairs, where the first element represents the operation and the second corresponds to the associated pulse shape. Using this training set, we can proceed to train a neural network.
@@
However, when we applied this approach to a three-qubit NMR system, the results were sub-optimal, with error rates failing to decrease satisfactorily. To address this issue, we explored utilizing the GRAPE algorithm for constructing a more effective training set. In this approach, we randomly selected a variety of single-qubit gates and employed the GRAPE algorithm to determine optimal control pulses for their implementation. A crucial aspect of our approach in using the GRAPE algorithm to train the network and achieve optimal results was ensuring that the optimization search started from a common initial point. Upon calculating the cosine similarity between the randomly generated pulses in our initial approach, we found that the similarity was relatively low. By instead using a common starting point in the GRAPE algorithm to develop pulses for the desired gates, we effectively constrained the search to a specific region within the space of control parameters. Consequently, within this region, the pulses exhibit a higher degree of similarity, as illustrated in Fig. 1.
@@
As we will demonstrate, employing this strategy significantly improves the cosine similarity between the resulting pulses. This approach helps us to focus on data around a specific point in the solution space. Intuitively, initializing the GRAPE algorithm from a random starting point for each operation leads to convergence toward very different pulse shapes. In contrast, using the same initial point for all operations results in similar pulse shapes across them. In a problem with multiple, or even infinitely many, local minima, it is crucial to limit our analysis of the dynamic system to a specific local solution. Ignoring this necessity results in severely non-smooth data geometry, preventing a simple, feasible neural network from finding any patterns within the feature space.
@@
Therefore, we have developed a neural network that, upon receiving a desired single-qubit gate as input, generates the corresponding pulse shape. Another key advantage of this approach is its ability to execute complicated operations within a single pulse. This method offers a significant improvement over sequential gate execution, particularly in the NISQ era, where accumulated gate fidelities can lead to a reduction in overall fidelity. By directly synthesizing the operation, we achieve higher fidelity with fewer steps.
@@ II Hamiltonian of the system and form of the control pulses
For an $n$-qubit NMR system with a homo-nuclear molecule, one can demonstrate that if the system is in the liquid regime at room temperature and operates with only one RF channel, then by selecting a rotating frame that matches the RF frequency of the channel, the total Hamiltonian in terms of the free evolution term and control term can be expressed as follow [42, 2, 3, 4]:
$H_{\text{total}}=H_{\text{free}}+H_{\text{ctrl}},$ (1)
@@
$H_{\text{ctrl}}=\sum_{i=1}^{n}\omega_{1}(t)\left(\sigma_{x}^{(i)}\cos\phi+\sigma_{y}^{(i)}\sin\phi\right).$ (3)
Here, $\omega_{0,i}$ represents the Larmor frequency of qubit $i$ under the static magnetic field, while $\omega_{1}(t)$ denotes the time-dependent Larmor frequency of the qubits under the applied RF field, both arising from the Zeeman interaction and given by:
$\omega_{0,i}=\gamma_{i}B_{0},\omega_{1}(t)=\gamma_{i}B_{1}(t),$ (4)
where $B_{0}$ is the static magnetic field, and $B_{1}(t)$ and $\phi(t)$ represent the amplitude and phase of the applied RF field, respectively. $\gamma_{i}$ is the gyromagnetic ratio for each nucleus in the system. Since $B_{1}(t)$ is significantly smaller than $B_{0}$, and the system is homonuclear, we use the approximation $\gamma_{i}B_{1}(t)=\omega_{1}(t)$ for any ii. The second term in $H_{\text{free}}$ represents the J-coupling between different nuclei [3].
To implement the desired quantum logic gates, we must determine $B_{1}(t)$ and $\phi(t)$. As a common technique in the GRAPE algorithm, these pulses are approximated as constant within NN discrete time steps. Within each step, the amplitude and phase remain constant, allowing us to compute the system’s overall operation by multiplying the operations from each time step. Since the Hamiltonian is constant within each step, we can readily calculate the dynamics for that period. Consequently, the total operation resulting from the pulse sequence is obtained by the product of the dynamics across all steps [43],
$U_{total}=u_{1}u_{2}.....u_{N-1}u_{N}$ (5)
where $u_{i}$ represents the unitary operation applied in step ii. In this algorithm, the derivative of fidelity with respect to control parameters is computed. By iteratively updating these parameters and following the gradient direction, we can converge on the optimal solution [15]. Here, we focus solely on implementing a single-qubit gate. The objective is for the overall system dynamics to produce the desired operation on the target qubit, while keeping all other qubits unaffected—i.e., they should undergo no evolution, that is:
$U_{target}=U_{1}\otimes I_{2}\otimes I_{3},$ (6)
where $U_{1}$ represents the effective evolution of qubit 1 after NN steps, and $I_{j}$ describes the effective evolution of qubit $j$, ensuring that its state returns to the initial state—i.e., no evolution occurs for that qubit. Our demand is that $U_{total}$ approaches the $U_{target}$ as closely as possible.
To simplify the learning process and reduce the number of control parameters, we used only phases to develop control pulses while keeping the amplitude constant. As we will see (and discussed in [40, 41]), this constraint does not significantly limit us in achieving acceptable fidelities, and we can reach high fidelities by only modulating the phase. To test the stability of this approach, after calculating the phase-modulated pulse with constant amplitude, we used it as the initial state for the GRAPE algorithm, by adding the amplitudes as an additional control parameters. The final result was not significantly different from the initial state, indicating that these pulses approximately satisfy the condition of zero gradient Euclidean norm with respect to phase elements, even after additional degrees of freedom related to amplitudes are added to the optimization problem.
@@ III Designing the neural network
As previously discussed, efficiently training the target neural network requires a set of sample pulses with high cosine similarity, ensuring that the pulses remain closely related. Cosine similarity quantifies the resemblance between two vectors by measuring the cosine of the angle between them in a multi-dimensional space. A value closer to 1 signifies a higher degree of similarity, indicating that the vectors—here, the control pulses—are structurally more alike. For efficient training, we focus on generating phase-only control pulses [40, 41]. In this approach, each pulse is represented as a vector, with its elements corresponding to the phase at specific time intervals. In this way, cosine similarity serves as a key metric for evaluating the similarity between sample pulses, enabling precise assessment and aiding in pulse selection for network training.
@@
A straightforward strategy for training the network involves generating a series of random pulses and computing the resulting operations they induce in the system. However, upon applying Algorithm 1 to compute the cosine similarity between these generated pulses, we observed that their similarity was not particularly significant. This highlights the need for a more refined approach to pulse selection to enhance network performance.
Algorithm 1 Cosine Similarity Analysis for the Control Pulses
1:for $i\leftarrow 1$ to $N_{\text{total}}$ do
2:  for $j\leftarrow 1$ to $N_{\text{total}}$ do
3:   $c[i,j]\leftarrow\dfrac{\langle\phi_{\text{data}}[i],\,\phi_{\text{data}}[j]\rangle}{\|\phi_{\text{data}}[i]\|\cdot\|\phi_{\text{data}}[j]\|}$ $\triangleright$ Compute the cosine similarity between pulse sequences of sample $i$ and $j$.
4:  end for
5:end for
We then consider utilizing the GRAPE algorithm to develop a training set for the neural network. Hereafter, we restrict ourselves to a three-qubit system and train the network to generate pulses for implementing any desired gate on qubit 1. Training the network to generate pulses for arbitrary gates on other qubits (i.e., qubits 2 and 3) follows the same procedure.
Algorithm 2 Input Data Generation Using Axis–Angle Representation with Cosine–Inverse Sampling
1:for $i\leftarrow 1$ to $N_{\text{total}}$ do
2:  Generate a uniformly random $2\times 2$ unitary matrix $U_{1}\in SU(2)$
3:  Sample random variables:
$\phi\sim\mathcal{U}(0,2\pi),\quad r\sim\mathcal{U}(0,1),\quad\gamma\sim\mathcal{U}(0,2\pi)$
4:  Compute the polar angle using the cosine–inverse approach:
$\theta=\arccos(1-2r)$
5:  Define the unit vector representing the rotation axis on the Bloch sphere:
$\hat{n}(\theta,\phi)=\begin{pmatrix}\sin\theta\cos\phi\\ \sin\theta\sin\phi\\ \cos\theta\end{pmatrix}$
6:  Construct the unitary matrix using the axis–angle representation:
$U_{1}=\cos\!\left(\frac{\gamma}{2}\right)I_{2}-i\sin\!\left(\frac{\gamma}{2}\right)\left(n_{x}\sigma_{x}+n_{y}\sigma_{y}+n_{z}\sigma_{z}\right)$
7:  Embed the $2\times 2$ unitary into an $8\times 8$ unitary matrix via Kronecker products
$U\leftarrow\mathrm{kron}(\mathrm{kron}(U_{1},I_{2}),I_{2})$
8:  Store real and imaginary parts
$X[i,:,:,1]\leftarrow\Re(U),\quad X[i,:,:,2]\leftarrow\Im(U)$
9:end for
@@
In first scenario, we used the GRAPE algorithm to obtain the optimal control pulses for implementing a typical single-qubit gate (Hadamard gate), each initialized from a different starting point. Figure 2 illustrates the cosine similarity of 16 samples generated using the GRAPE algorithm for a specific logic gate, each initialized with a different starting point. The similarity among pulses generated by the GRAPE algorithm for the same gate but with different starting points is relatively low.
In the second scenario, we selected 17,000 uniformly chosen single-qubit gates (Algorithm 2) and applied the GRAPE algorithm to determine their optimal control pulses, maintaining a common starting point across all instances. To construct this uniform set, we adopt the axis-angle representation of $U(2\times 2)$ unitary matrices in the Bloch sphere, defined as $U(\theta,\phi,\gamma)=e^{-i\sigma\cdot\hat{n}(\theta,\phi)\gamma/2}$, where $\hat{n}(\theta,\phi)$ represents the rotation axis and $\gamma$ denotes the rotation angle in the Bloch sphere representation. By uniformly selecting $\hat{n}(\theta,\phi)$ on the sphere using the cosine inverse approach and also uniformly selecting the rotation angle $\gamma$ around the axis we construct the desired set of uniform unitary matrices. Surprisingly, as shown in Fig. 3, the cosine similarity analysis of these 17,000 samples reveals a high degree of similarity among the pulses, despite their correspondence to different gates. A comparison between Figures 2 and 3 reveals a noticeable increase in similarity among the samples when the GRAPE algorithm is used with a common starting point. This highlights the importance of maintaining a fixed initialization when generating training pulses, as it effectively constrains the search to a specific region within the space of control parameters, ensuring that the resulting pulses remain closely related.
@@
Figure 2: Cosine similarity analysis of 16 samples generated using the GRAPE algorithm for a specific logic gate, each with a different starting point.The horizontal (x) and vertical (y) axes represent the indices of sample pulses. Value of the similarity between the pulses x and y is shown on the x-y cells. The diagonal entries all equal 1, as the cosine similarity between a pulse and itself is always 1.
Figure 3: Cosine similarity analysis of 17,000 samples generated using the GRAPE algorithm for the random logic gates, each with a same starting point.The horizontal (x) and vertical (y) axes represent the indices of sample pulses. Value of the similarity between the pulses x and y is shown on the x-y cells.
Therefore, we adopted the second scenario to generate sample pulses corresponding to selected unitary operations, forming the ground truth for training data-set, ensuring efficient network training (Algorithm 3).
Algorithm 3 Output Ground Truth for Training Set
1:for all Unitary $i$ in X do
2:  $G_{\text{target}}\leftarrowX$[i,:,:,1]+j.X[i,:,:,2]X[i,:,:,1]+j.X[i,:,:,2], where j is the imaginary unit.
3:  Initialize control parameters: $A\leftarrow A_{0}$ and $\phi\leftarrow\phi_{0}=\begin{bmatrix}\pi&\pi&\pi&\dots&\pi\end{bmatrix}$ , where $A_{0}$ is assumed to be uniform on time.
4:  for $iter\leftarrow 1$ to num_iterationsnum\_iterations do
5:   $d\phi\leftarrow\textsc{Deri\_func}(A_{0},\phi,G_{\text{target}})$ $\triangleright$ Compute gradients of the fidelity cost with respect to $\phi$, using the methodology in [4]. Treating $A_{0}$ not to be a part of the optimization set.
6:   Update $\phi$ using $d\phi$ and ADAM Gradient Ascent [44].
7:  end for
8:  Save optimized pulse parameters: $\phi_{\text{data}}[i]\leftarrow\phi$.
9:end for
After creating 17,000 samples, we trained a feed-forward neural network [45] to establish the mapping function between the overall operation applied to the system and control pulses (Algorithm 4). The trained neural network is shown in Fig. 4. To implement the layers of the neural network, we utilized the architecture depicted in Fig. 5.
Algorithm 4 Neural Network Design and Training for Pulse Generation
1:Inp $\leftarrow$ For each generated gate, U (the 2×2 matrix), flatten and concatenated the real and imaginary components, into a one-dimensional input array with 8 elements.
2:Create the complete dataset structure $(Inp,\phi_{\text{data}})$, and split the dataset into training set and test set.
3:Define a deep feed-forward neural network with:
  • Input layer: 8 features.
  • Hidden layers: Several fully-connected layers with non-linear activations (e.g., tanh, ReLU) and dropout layers.
  • Output layer: $T$ real-valued outputs corresponding to the predicted pulse phases, T is the length of the phase pulse, $\phi$.
4:Set hyperparameters and use back propagation to train the network. $\triangleright$ Repeat by using various architectures and combinations of hyperparameters, until specific convergence criteria are met. Regulate the dropout rate and L2 norm to prevent overfitting. $\triangleright$ The network learns the mapping from unitary features to optimal pulse parameters.
Figure 4: The neural network receives a $U$ (2×2 unitary matrix), representing the gate to be applied to the target qubit (for example, qubit 1). It flattens and concatenates the real and imaginary components into a one-dimensional input array with 8 elements. The network then interprets the overall evolution as $U_{1}\otimes I_{2}\otimes I_{3}$ and generates the pulse required to implement the desired gate, which is transmitted to the output. Upon applying the pulse, the operator $U_{f}$ is executed on the system.
Figure 5: Neural Network Architecture Overview. This diagram presents the structured flow of the neural network, beginning with the feature input and advancing through sequential layers that include fully connected layers, dropout modules, and activation functions (ReLU and Tanh). The final stage culminates in a regression output, illustrating how the network processes and transforms input data to generate predictions.